### v1

# Fundamental Limits of Coded Polynomial Aggregation

Abstract
We introduce the problem of coded polynomial aggregation (CPA) in distributed computing systems, where the goal is to recover a weighted sum of polynomial evaluations using responses from distributed workers. A fundamental question in CPA is the minimum number of responses required to recover the computation. A straightforward approach is individual decoding, where the system first decodes all individual polynomial evaluations and then aggregates them, which requires at least d(K1)+1d(K-1)+1 responses for a degree-$d$ polynomial evaluated on $K$ data points. In this paper, we establish a necessary and sufficient condition for exact recovery when the number of responses satisfies $N\leq d(K-1)$, a regime where individual decoding is infeasible. Moreover, we characterize the minimum number of responses required by CPA, showing that it equals $\left\lfloor\frac{K-1}{2}\right\rfloor+1$ when $d=1$, and (d1)(K1)+1(d-1)(K-1)+1 when $d\geq 2$. These thresholds are strictly smaller than those required by a straightforward strategy based on individual decoding. We further provide explicit constructions of CPA schemes that achieve this bound.
I Introduction
Coded distributed computing is a fundamental paradigm for executing large-scale data processing tasks, where a master node partitions a computation task, assigns sub-tasks to multiple workers, and aggregates their responses to recover the desired computation. One commonly used metric in coded distributed computing systems is the minimum amount of information that must be collected from the workers to recover the computation. This metric can be characterized by the minimum number of responses, defined as the minimum number of worker responses required for successful decoding.
For matrix multiplications and polynomial computations, many prior works apply polynomial-based coded computing techniques. Specifically, the authors in [5] use one-dimensional maximum distance separable (MDS) codes to encode matrices, achieving a minimum number of responses equal to the partition level of the data matrix. Product codes [6] use two-dimensional MDS codes and apply iterative decoding algorithms. Polynomial codes [14] encode sub-matrices as coefficients of a univariate polynomial and interpolate this polynomial using responses from workers. PolyDot codes and MatDot codes [1] extend the interpolation-based approach to trade off computation and communication costs by partitioning datasets. Entangled polynomial codes [12, 15] exploit bilinear structure and extend polynomial coded computing to scenarios such as secure, private, and batch computation. To construct numerically stable interpolation schemes, the authors in [2] apply Chebyshev polynomials to coded computing. Instead of univariate polynomials, bivariate polynomial coded schemes [3] develop two-dimensional interpolation, enabling finer trade-offs among storage, computation, and communication. Differing from encoding sub-computations as coefficients of a polynomial, Lagrange coded computing [13] encodes sub-computations as evaluations of Lagrange polynomials, achieving lower decoding complexity.
Apart from decoding exact computations, several works consider approximate computation of arbitrary functions. The Berrut-approximated coded computing scheme proposed in [4] replaces polynomial interpolation with rational interpolation to enable numerically stable approximation. In 2025, the authors in [8] proposed learning-theoretic coded computing, by formulating coded computing as an encoder–decoder design problem that minimizes the mean squared error (MSE), and derived optimal coding schemes based on smoothing splines. This approach is extended to the scenario where each worker becomes a straggler with some probability [7].
A common paradigm underlying many existing coded computing schemes is individual decoding, where the master reconstructs all individual sub-computations before computing the desired output. In particular, for matrix multiplications, the works in [5, 6, 14, 1, 12, 15, 2, 3] decode all sub-matrices contained in the desired matrix, and therefore recover the computation. However, for aggregation tasks, individual decoding is sufficient but not necessary, and may lead to redundant computation. To compute aggregation tasks in the form of gradient computations, [11] proposed gradient coding, which introduces redundancy into gradient computations to mitigate slow or unavailable workers. A different class of aggregation schemes considers linearly separable computations [10], where the authors focus on computing weighted aggregations of arbitrary computation functions while minimizing communication under uncoded data storage.
Despite these extensive efforts, the problem of computing a weighted aggregation of polynomial computations has not been studied as a standalone objective. Specifically, we consider the computation of weighted aggregation,
$\bm{Y}\triangleq\sum_{k=0}^{K-1}w_{k}\,F(\bm{X}_{k}),$ (1)
where $w_{k}\in\mathbb{C}$ are non-zero weights, and $F(\cdot)$ is a polynomial of degree $d$ that operates element-wise on each data matrix $\bm{X}_{k}\in\mathbb{C}^{q\times v}$. This formulation commonly arises in distributed learning and signal processing, where the objective is to recover an aggregated output rather than all individual computations. A natural baseline in coded polynomial computing is individual decoding, where workers compute polynomial evaluations on encoded data, and the master interpolates the resulting polynomial to recover all $F(\bm{X}_{k})$ for $k\in[K]$. Under individual decoding, the required number of responses for exact recovery is d(K1)+1d(K-1)+1. However, the objective is to recover a weighted aggregation rather than each $F(\bm{X}_{k})$, which leads to the following questions.
  1. 1.
    Can the structures of aggregation and polynomials be exploited to reduce the number of worker responses for exact recovery?
  2. 2.
    What is the minimum number of responses?
  3. 3.
    Do there exist explicit schemes that achieve this minimum number of responses?
In this paper, we address these questions by characterizing the fundamental limits of coded polynomial aggregation (CPA) and constructing optimal schemes, as summarized below.
  1. 1.
    We propose a framework of CPA to compute the weighted aggregation in (1), and establish a necessary and sufficient condition for exact recovery in the regime where individual decoding is infeasible.
  2. 2.
    We characterize the minimum number of responses $N^{*}$ required for a feasible CPA scheme.
  3. 3.
    We provide explicit schemes that achieve this bound $N^{*}$.
II Problem Formulation
We consider a distributed computing system consisting of a master node and a set of $N$ worker nodes, indexed by $N]\triangleq\{0,1,\ldots,N-1\}$. Given $K$ data matrices $\bm{X}_{k}\in\mathbb{C}^{q\times v}$ for $k\in[K]$, a polynomial function $F(\cdot)$ of degree $d$, and a weight vector $\bm{w}\in\mathbb{C}^{K}$, where each element $w_{k}\neq 0$ for all $k\in[K]$, the objective of the master is to compute the weighted aggregation defined in (1), using the responses from workers. Each worker returns one response to the master.
II-A Proposed CPA
A CPA scheme consists of three phases: encoding, computing, and decoding.
II-A1 Encoding
The master selects a set of $K$ distinct numbers $\{\alpha_{k}$ $\in$ $\mathbb{C}:$ $k\in[K]\}$, referred to as the data points, and interpolates an encoder polynomial $E(z)$ such that $E(\alpha_{k})$ == $\bm{X}_{k}$ for all $k\in[K]$. Next, the master selects another set of $N$ distinct numbers $\{\beta_{n}\in\mathbb{C}:$ $n\in[N]\}$, referred to as the evaluation points, satisfying $\{\alpha_{k}:$ $k\in[K]\}$ $\cap$ $\{\beta_{n}:$ $n\in[N]\}$ == $\emptyset$. The master evaluates $E(z)$ at $\{\beta_{n}:n\in[N]\}$ and sends the coded matrix $E(\beta_{n})$ to worker $n$.
II-A2 Computing
Each worker $n\in[N]$ computes $F(E(\beta_{n}))$ locally and returns the result to the master.
II-A3 Decoding
Upon receiving the responses from all $N$ workers, the master interpolates a decoder polynomial $D(z)$ such that $D(\beta_{n})=F(E(\beta_{n}))$ for all $n\in[N]$. The master computes $D(\alpha_{k})$ for $k\in[K]$, and obtains
$\widehat{\bm{Y}}\triangleq\sum_{k=0}^{K-1}w_{k}\,D(\alpha_{k}).$
We define the feasibility of a CPA scheme as follows.
Definition 1 (Feasibility of CPA)
Fix positive integers $K$, $d$, and $N$, data points $\{\alpha_{k}$ :: $k\in[K]\}$ and evaluation points $\{\beta_{n}:n\in[N]\}$. A CPA scheme is feasible if it produces $\widehat{\bm{Y}}=\bm{Y}$. $\diamond$
We define the minimum number of responses as follows.
Definition 2 (Minimum Number of Responses)
For positive integers $K$ and $d$, and a given set of data points $\{\alpha_{k}:k\in[K]\}$, the minimum number of responses is defined as the smallest integer $N$ for which there exist evaluation points $\{\beta_{n}:n\in[N]\}$ such that a feasible CPA scheme exists. $\diamond$
In this paper, the goal is to characterize the minimum number of responses for which a feasible CPA scheme exists, and to construct explicit schemes that achieve this bound.
II-B Individual Decoding Baseline
Existing results on polynomial coded computing, including [14, 1, 12, 15, 2, 3, 13], recover the desired computation by decoding all individual sub-computations through reconstructing polynomial evaluations or coefficients. Among these works, we set Lagrange coded computing [13] as the baseline to our setting. It focuses on polynomial computations and encodes each data matrix $\bm{X}_{k}$ as an evaluation of an encoder polynomial $E(z)$. When applied to the CPA setting, Lagrange coded computing leads to the following CPA based on individual decoding.
Definition 3 (CPA Based on Individual Decoding)
We define CPA based on individual decoding as a decoding strategy, where the master reconstructs all individual evaluations $F(\bm{X}_{k})$ by interpolating a polynomial $D(z)$ satisfying $F(\bm{X}_{k})=D(\alpha_{k})$ for all $k\in[K]$. $\diamond$
Figure 1: Feasible regimes of CPA and the baseline CPA scheme based on individual decoding for computation degrees $d=1$ and $d=2$. For a fixed $K$, the feasible regime consists of the values of $N$ for which the corresponding scheme is feasible. The red region corresponds to CPA, which enables feasibility with fewer responses ($N\leq d(K-1)$) via the orthogonality conditions in Theorem 1, with the minimum number of responses $N^{\star}$ characterized in Theorem 2. The gray region corresponds to the baseline scheme based on individual decoding, which is feasible only when $N\geq d(K-1)+1$, as stated in Lemma 1.
The following lemma characterizes the minimum number of responses for a CPA scheme based on individual decoding.
Lemma 1
For integers $K$ and $d$, a CPA scheme based on individual decoding is feasible if and only if $N\geq d(K-1)+1$.
Proof:
The master recovers all individual evaluations by interpolating the composed polynomial $D(z)=F(E(z))$, so that $D(\alpha_{k})=F(\bm{X}_{k})=F(E(\alpha_{k}))$ for all $k\in[K]$. Since $\deg(E)\leq K-1$ and $F(z)$ has degree $d$, we have $\deg(D)\leq d(K-1)$. Therefore, $D(z)$ can be uniquely interpolated from the worker responses if and only if $N\geq d(K-1)+1$. ∎
Lemma 1 shows that CPA schemes based on individual decoding require at least d(K1)+1d(K-1)+1 responses.
In this paper, we study CPA in the regime $N\leq d(K-1)$, where individual decoding is infeasible. We show that exact recovery remains possible with strictly fewer worker responses by directly exploiting the aggregation structure.
III Main Results
III-A Necessary and Sufficient Orthogonality Conditions
We characterize a necessary and sufficient condition for the existence of a feasible CPA scheme when $N\leq d(K-1)$.
Theorem 1
For positive integers $K$, $d$, and $N$ with $2\leq N\leq d(K-1)$, let $C\triangleq d(K-1)-N+1$. A CPA scheme is feasible if and only if the data points $\{\alpha_{k}\in\mathbb{C}:k\in[K]\}$ and evaluation points $\{\beta_{n}\in\mathbb{C}:n\in[N]\}$ satisfy $\{\alpha_{k}:k\in[K]\}\cap\{\beta_{n}:n\in[N]\}=\emptyset$, and
$\sum_{k=0}^{K-1}w_{k}\,P(\alpha_{k})\,\alpha_{k}^{j}=0,\quad\forall j\in[C],$ (2)
where
$P(z)\triangleq\prod_{n\in[N]}(z-\beta_{n}).$ (3)
Proof:
The proof is provided in Section IV. ∎
Theorem 1 shows that a CPA scheme is feasible if and only if the data points $\{\alpha_{k}:k\in[K]\}$ and evaluation points $\{\beta_{n}:n\in[N]\}$ satisfy a system of orthogonality conditions, as shown in (2). The intuition behind the proposed conditions in (2) is as follows. When $N\leq d(K-1)$, the decoder polynomial $D(z)$ has a strictly smaller degree than the polynomial $F(E(z))$, and therefore cannot recover the individual evaluations $F(E(\alpha_{k}))$. However, the recovery error $\widehat{\bm{Y}}-\bm{Y}$ can be expressed as a linear combination of $\sum_{k=0}^{K-1}w_{k}P(\alpha_{k})\alpha_{k}^{j}$ for $j\in[C]$. The orthogonality conditions in (2) force all such quantities to be zero, thereby eliminating the recovery error and enabling exact recovery of $\bm{Y}$.
III-B Minimum Number of Responses
The following theorem characterizes the minimum number of responses required for the existence of a feasible CPA scheme for a given generic set of pairwise distinct data points $\{\alpha_{k}:k\in[K]\}$, where the data points are chosen outside a proper algebraic variety depending on the system parameters $K$, NN, and $d$, and satisfy $\alpha_{i}\neq\alpha_{j}$ for all $i\neq j$.
Theorem 2
For positive integers $K$, $d$, and $N$ with $2\leq N\leq d(K-1)$, and for a given generic set of pairwise distinct data points $\{\alpha_{k}:k\in[K]\}$, there exists evaluation points $\{\beta_{n}:n\in[N]\}$ such that CPA is feasible if and only if $N\geq N^{*}$, where
$N^{*}=\begin{cases}\left\lfloor\dfrac{K-1}{2}\right\rfloor+1,&\text{if }d=1,\\[6.0pt] (d-1)(K-1)+1,&\text{if }d\geq 2.\end{cases}$ (4)
From Theorem 2, when $d=1$, the minimum number of responses required by CPA scales to approximately half of that required by individual decoding as $K$ grows large. When $d\geq 2$, CPA reduces the required number of responses by K1K-1 compared to the individual decoding baseline in Lemma 1.
Fig. 1 illustrates the minimum number of responses required by CPA and the baseline CPA schemes based on individual decoding. As shown in Fig. 1, for both $d=1$ and $d=2$, CPA has a strictly smaller minimum number of responses for any $K$. Moreover, the difference between the proposed CPA schemes and the baseline CPA scheme based on individual decoding grows with $K$, indicating that the benefit of CPA becomes more pronounced as the number of data matrices increases.
III-C Achievable CPA Schemes
We next present an achievable CPA scheme by constructing evaluation points $\{\beta_{n}:n\in[N]\}$ for a given generic set of pairwise distinct data points $\{\alpha_{k}:k\in[K]\}$.
Theorem 3
For positive integers $K$, $d$, and NN satisfying $N^{*}\leq N\leq d(K-1)$, and for a given generic set of pairwise distinct data points $\{\alpha_{k}:k\in[K]\}$, Algorithm 1 constructs evaluation points $\{\beta_{n}:n\in[N]\}$ that satisfy the conditions in Theorem 1.
Proof:
The proof is provided in Section V. ∎
Algorithm 1 Construction of $\{\beta_{n}:n\in[N]\}$
1:$K,d,N$, $\bm{w}$, $\{\alpha_{k}:k\in[K]\}$
[…]
3:Construct $\bm{V}\in\mathbb{C}^{C\times K}$ with $V[j,k]=\alpha_{k}^{\,j}$ for $j\in[C]$, $k\in[K]$
4:Construct $\bm{A}\in\mathbb{C}^{K\times(N+1)}$ with $A[k,n]=\alpha_{k}^{\,n}$ for $k\in[K]$, $n\in[N+1]$
[…]
6:Select a nonzero vector $\bm{c}=(c_{0},\ldots,c_{N})^{\mathsf{T}}\in\ker(\bm{U})$
7:Form the polynomial $P_{\text{form}}(z)=\sum_{n=0}^{N}c_{n}z^{n}$
8:Compute the $N$ roots $\{\beta_{n}:n\in[N]\}$ of $P_{\text{form}}(z)$
9:return $\{\beta_{n}:n\in[N]\}$
The following example illustrates the construction in Algorithm 1 for the case $d=1$.
Example 1
Consider the linear case $d=1$ with $K=3$ and data points $\{\alpha_{0}$, $\alpha_{1}$, $\alpha_{2}\}$ == $\{-1$, 0, 1}1\}, and weights $\bm{w}=(-0.5$, $1$, $0.5)$. Under individual decoding, d(K1)d(K-1) +1+1 =3=3 workers are required, and the evaluation points $\{\beta_{0}$, $\beta_{1}$, $\beta_{2}\}$ can be chosen arbitrarily as long as they are distinct and disjoint from $\{-1$, 0, 1}1\}.
In contrast, CPA allows $N=2$, which gives $C=1$. Applying Algorithm 1, we obtain $\bm{V}=\begin{bmatrix}1&1&1\end{bmatrix}\text{ and }\bm{A}=\begin{bmatrix}1&-1&1\\ 1&0&0\\ 1&1&1\end{bmatrix}$, which yields $\bm{U}=\begin{bmatrix}1&1&0\end{bmatrix}$. Solving $\bm{U}\bm{c}=\bm{0}$ gives $\bm{c}$ =(0.0506=(-0.0506, $0.0506$, $0.5)$ and $P_{\text{form}}(z)$ =0.0506=-0.0506 +0.0506z+0.0506z $0.5z^{2}$. The two roots are $\beta_{0}=-0.37272$ and $\beta_{1}=0.27152$, which are distinct and disjoint from $\{-1,0,1\}$. Finally, the orthogonality conditions in Theorem 1 are numerically satisfied, with $\sum_{k=0}^{2}w_{k}P(\alpha_{k})\approx 5.5\times 10^{-17}$.
IV Proof of Theorem 1
We start by considering the scalar case where $\bm{X}_{k}$ reduces to a scalar $x_{k}$ and the desired output $\bm{Y}$ reduces to a scalar $y$, i.e., $q=v=1$. The extension to the matrix case follows directly by applying the same argument element-wise, since the polynomial $F(\cdot)$ operates element-wise on the data matrices. Define the error polynomial $\Delta(z)\triangleq D(z)-F(E(z))$. Its degree satisfies $\deg(\Delta)\leq\max\{\deg(D),\deg(F(E))\}\leq\max\{N-1,d(K-1)\}=d(K-1)$. The decoding error $\widehat{y}-y$ can be expressed by
$\widehat{y}-y=\sum_{k=0}^{K-1}w_{k}\Delta(\alpha_{k}),$ (5)
as $\widehat{y}-y$ == $\sum_{k=0}^{K-1}w_{k}\big(D(\alpha_{k})-F(x_{k})\big)$ $\overset{(a)}{=}$ $\sum_{k=0}^{K-1}w_{k}\big(D(\alpha_{k})-F(E(\alpha_{k}))\big)$ == $\sum_{k=0}^{K-1}w_{k}\Delta(\alpha_{k})$, where $(a)$ is due to $x_{k}=E(\alpha_{k})$. During decoding, $D(\beta_{n})=F(E(\beta_{n}))$ imply $\Delta(\beta_{n})$ == 0 for all $n\in[N]$. Consequently, we can write $\Delta(z)$ as
$\Delta(z)=\prod_{n\in[N]}(z-\beta_{n})\cdot R(z)=P(z)\cdot R(z),$ (6)
where $P(z)=\prod_{n\in[N]}(z-\beta_{n})$ has degree $N$, as shown in (3), and $R(z)$ is a polynomial satisfying $\deg(R)$ == $\deg(\Delta)-\deg(P)$ $\leq$ d(K1)Nd(K-1)-N. Writing $R(z)$ == $\sum_{j=0}^{\deg(R)}r_{j}z^{j}$, where $\{r_{j}:$ $j\in[\deg(R)+1]\}$ are arbitrary complex coefficients. We substitute (6) into (5) yields
$\widehat{y}-y$ $=\sum_{k=0}^{K-1}w_{k}P(\alpha_{k})R(\alpha_{k})$
$=\sum_{k=0}^{K-1}w_{k}P(\alpha_{k})\sum_{j=0}^{\deg(R)}r_{j}\alpha_{k}^{j}$
$=\sum_{j=0}^{\deg(R)}r_{j}\left(\sum_{k=0}^{K-1}w_{k}P(\alpha_{k})\alpha_{k}^{j}\right).$
Therefore, $\widehat{y}-y=0$ for all admissible choices of $R(z)$ if and only if $\sum_{k=0}^{K-1}w_{k}P(\alpha_{k})\alpha_{k}^{j}=0$ for all $j=$ 0, 11, $\ldots$, d(K1)Nd(K-1)-N. This establishes the necessary and sufficient condition for feasibility in the scalar case. The same argument applies element-wise to the matrix-valued case, and hence the condition remains necessary and sufficient for feasibility. This completes the proof of Theorem 1.
V Proof of Theorem 2 and Theorem 3
To prove Theorem 2, it is equivalent to show that there exists a set $\{\beta_{n}:n\in[N]\}$ satisfying the following three properties:
$\textbf{(Orthogonality): }\sum_{k=0}^{K-1}w_{k}\,P(\alpha_{k})\,\alpha_{k}^{j}=0,\quad j\in[C].$ (7)
$\textbf{(Disjointness): }\{\alpha_{k}:k\in[K]\}\cap\{\beta_{n}:n\in[N]\}=\emptyset.$ (8)
$\textbf{(Distinction): }\beta_{n}\neq\beta_{n^{\prime}}\text{ for all }n\neq n^{\prime}.$ (9)
if and only if $N\geq N^{*}$, where $N^{*}$ is given in (4).
V-A Reformulation of Orthogonality
Let $\operatorname{diag}(\bm{w})$ denote the diagonal matrix with diagonal entries given by $\bm{w}$. Define $\bm{p}=(p_{0},\ldots,p_{K-1})^{\mathsf{T}}\in\mathbb{C}^{K}$ with $p_{k}\triangleq P(\alpha_{k})$. Let $\bm{V}\in\mathbb{C}^{C\times K}$ be the Vandermonde matrix with entries $V[j,k]=\alpha_{k}^{j}$ for $j\in[C]$ and $k\in[K]$. Substituting $\bm{V}$ and $\bm{p}$ into (7) yields
$\sum_{k=0}^{K-1}w_{k}\,P(\alpha_{k})\,\alpha_{k}^{j}=0,j\in[C]\ \Leftrightarrow\ \bm{V}\,\operatorname{diag}(\bm{w})\,\bm{p}=\bm{0}.$ (10)
Next, we express $\bm{p}$ in terms of the coefficients of the polynomial $P(z)$. Let $P_{\mathrm{form}}(z)=c_{0}+c_{1}z+\ldots+c_{N}z^{N}$ be a polynomial with distinct roots $\{\beta_{n}:n\in[N]\}$, where $c_{N}\neq 0$. Equivalently, $P_{\mathrm{form}}(z)=c_{N}(z-\beta_{0})(z-\beta_{1})\ldots(z-\beta_{N-1})$. From (3), $P(z)=\frac{1}{c_{N}}P_{\mathrm{form}}(z)$ for some coefficient vector $\bm{c}=(c_{0},\ldots,c_{N})^{\mathsf{T}}\in\mathbb{C}^{N+1}$ with $c_{N}\neq 0$. Then,
$\bm{p}=\begin{bmatrix}P(\alpha_{0})&P(\alpha_{1})&\ldots&P(\alpha_{K-1})\end{bmatrix}^{\mathsf{T}}=\frac{1}{c_{N}}\,\bm{A}\bm{c},$ (11)
where we define $\bm{A}\in\mathbb{C}^{K\times(N+1)}$ with entries $A[k,n]=\alpha_{k}^{n}$ for $k\in[K]$ and $n\in[N+1]$. Substituting (11) into (10) yields
$\bm{V}\operatorname{diag}(\bm{w})\bm{p}=\bm{0}$ $\overset{(a)}{\Leftrightarrow}\bm{V}\operatorname{diag}(\bm{w})\bm{A}\bm{c}\cdot\frac{1}{c_{N}}=\bm{0}$ (12)
$\overset{(b)}{\Leftrightarrow}\bm{V}\operatorname{diag}(\bm{w})\bm{A}\bm{c}=\bm{0}$
$\overset{(c)}{\Leftrightarrow}\bm{U}\bm{c}=\bm{0},$
where $(a)$ follows from (11), $(b)$ follows from $c_{N}\neq 0$, and in $(c)$ we define $\bm{U}\triangleq\bm{V}\operatorname{diag}(\bm{w})\bm{A}\in\mathbb{C}^{C\times(N+1)}$.
In $\bm{U}\bm{c}=\bm{0}$, the vector $\bm{c}$ corresponds to the coefficients of the polynomial whose roots are $\{\beta_{n}:n\in[N]\}$. The matrix $\bm{U}$ depends only on $\bm{w}$ and $\{\alpha_{k}:k\in[K]\}$ and satisfies
$\operatorname{rank}(\bm{U})=\min(C,K,N+1),$ (13)
as $\operatorname{rank}(\bm{U})$ == $\operatorname{rank}(\bm{V}\operatorname{diag}(\bm{w})\bm{A})$ $\leq$ $\min(\operatorname{rank}(\bm{V}$ $\operatorname{diag}(\bm{w}))$, $\operatorname{rank}(\bm{A}))$ $\overset{(a)}{=}$ $\min(\operatorname{rank}(\bm{V})$, $\operatorname{rank}(\bm{A}))$ == $\min(C$, $K$, $N+1)$, where $(a)$ follows from the fact that $\operatorname{diag}(\bm{w})$ is invertible.
V-B Necessity of Theorem 2
Assume that there exist distinct data points $\{\alpha_{k}:k\in[K]\}$ satisfying the three properties in (7)–(9).
V-B1 $C<K$ is necessary
Assume $C\geq K$. Then $\operatorname{rank}(\bm{V}\operatorname{diag}(\bm{w}))$ == $\operatorname{rank}(\bm{V})$ == $\min(C$, $K)$ =K=K. From (10), $\operatorname{dim}(\operatorname{ker}(\bm{V}\operatorname{diag}(\bm{w})))=K-\operatorname{rank}(\bm{V}\operatorname{diag}(\bm{w}))=0$, which implies $\bm{p}=\bm{0}$, i.e., $P(\alpha_{k})=0$ for all $k\in[K]$. This forces $\{\alpha_{k}:k\in[K]\}\subseteq\{\beta_{n}:n\in[N]\}$, contradicting the disjointness condition in (8). Therefore,
$C<K\Leftrightarrow N\geq(d-1)(K-1)+1.$ (14)
V-B2 $C<N+1$ is necessary
Since there exist non-trivial solutions to $\bm{U}\bm{c}=\bm{0}$ with $\bm{c}\neq\bm{0}$, we must have $\operatorname{dim}(\operatorname{ker}(\bm{U}))>0$. Moreover, $\operatorname{dim}(\operatorname{ker}(\bm{U}))=(N+1)-\operatorname{rank}(\bm{U})\geq(N+1)-\min(C,K,N+1)=(N+1)-\min(C,N+1)$. Hence, $(N+1)-\min(C,N+1)>0$, which implies
$N+1>C\Leftrightarrow N\geq\left\lfloor\frac{d(K-1)}{2}\right\rfloor+1.$ (15)
Combining (14) and (15), we obtain NN $\geq$ $\max((d-1)(K-1)+1$, $\left\lfloor\frac{d}{2}(K-1)\right\rfloor$ +1)+1), which yields (4). This completes the proof of necessity.
V-C Sufficiency of Theorem 2
Assume $N\geq N^{*}$. We construct a set of points $\{\beta_{n}:n\in[N]\}$ satisfying the conditions in (7)-(9). The key steps of the construction are as follows. 1) We solve the linear system $\bm{U}\bm{c}=\bm{0}$. 2) We select a vector $\bm{c}\in\ker(\bm{U})$ with $c_{N}\neq 0$ and obtain the polynomial $P(z)=\frac{c_{0}}{c_{N}}+\frac{c_{1}}{c_{N}}z+\ldots+z^{N}$. From (12), the orthogonality condition in (7) is satisfied. 3) We compute the $N$ roots of the polynomial $P(z)$, i.e., $\{\beta_{n}:n\in[N]\}$.
Importantly, we will show that for a generic set of pairwise distinct $\{\alpha_{k}:k\in[K]\}$, all but a Lebesgue measure-zero subset of $\bm{c}\in\ker(\bm{U})$ lead to evaluation points $\{\beta_{n}:n\in[N]\}$ satisfying (8) and (9). For background on relevant concepts in algebraic geometry, such as affine varieties and Zariski-open dense subsets, we refer the reader to [9].
V-C1 Proof of Orthogonality
Under $N\geq N^{*}$, we have $\operatorname{dim}(\operatorname{ker}(\bm{U}))>0$. Therefore, there exists a nonzero vector $\bm{c}\in\ker(\bm{U})$ such that $\bm{U}\bm{c}=\bm{0}$, as shown in equation $(c)$ of (12). Equation $(b)$ in (12) then implies that the polynomial $P_{\mathrm{form}}(z)=c_{0}+c_{1}z+c_{2}z^{2}+\ldots+c_{N}z^{N}$ satisfies the orthogonality condition $\sum_{k=0}^{K-1}w_{k}\,P_{\mathrm{form}}(\alpha_{k})\,\alpha_{k}^{j}=0$ for all $j\in[C]$. To ensure that equation $(a)$ in (12) holds, i.e., that the normalized polynomial $P(z)=\frac{P_{\mathrm{form}}(z)}{c_{N}}$ satisfies the orthogonality condition in (7), it is required that $c_{N}\neq 0$. This requirement is captured by the following lemma.
Lemma 2
Define a Zariski open subset of $\ker(\bm{U})$ as $\mathcal{C}_{N}\triangleq\{\bm{c}\in\ker(\bm{U}):c_{N}\neq 0\}$. For generic $\{\alpha_{k}:k\in[K]\}$, $\mathcal{C}_{N}\neq\emptyset$.
Proof:
The proof is sketched as follows. The set of $\bm{\alpha}=(\alpha_{0}$, $\ldots$ , $\alpha_{K-1})$ $\in$ $\mathbb{C}^{K}$ for which $\mathcal{C}_{N}$ == $\emptyset$ defines a proper affine variety in $\mathbb{C}^{K}$, and therefore has Lebesgue measure zero. Hence, for generic $\{\alpha_{k}:k\in[K]\}$, we have $\mathcal{C}_{N}\neq\emptyset$. The details are provided in Appendix A. ∎
From Lemma 2, for generic $\{\alpha_{k}:k\in[K]\}$, there exists a vector $\bm{c}\in\ker(\bm{U})$ satisfying $c_{N}\neq 0$. For such a choice of $\bm{c}$, the polynomial $P(z)=\frac{c_{0}}{c_{N}}+\frac{c_{1}}{c_{N}}z+\frac{c_{2}}{c_{N}}z^{2}+\ldots+z^{N}$ satisfies (7), and its NN roots are the desired $\{\beta_{n}:n\in[N]\}$.
V-C2 Proof of Disjointness
Condition (8) is equivalent to requiring $P(\alpha_{k})\neq 0$ for all $k\in[K]$. Since $P(z)$ and $P_{\mathrm{form}}(z)$ have the same roots, it suffices to require $P_{\mathrm{form}}(\alpha_{k})\neq 0$ for all $k\in[K]$. From (11), i.e., $P_{\mathrm{form}}(\alpha_{k})=[\bm{A}\bm{c}]_{k}$, it suffices to show the existence of a vector $\bm{c}\in\ker(\bm{U})$ such that $\bm{A}\bm{c}]_{k}\neq 0$ for all $k\in[K]$. This requirement is captured by the following lemma.
Lemma 3
Define a Zariski open subsets of $\ker(\bm{U})$ as $\mathcal{A}_{k}$ $\triangleq$ $\{\bm{c}\in\ker(\bm{U}):$ $\bm{A}\bm{c}]_{k} \neq$ 0}0\}, and let $\mathcal{C}_{A}$ $\triangleq$ $\bigcap_{k=0}^{K-1}$ $\mathcal{A}_{k}$. For generic $\{\alpha_{k}:k\in[K]\}$, the set $\mathcal{C}_{A}$ is nonempty.
Proof:
The proof is sketched as follows. Fix $k\in[K]$. The set of parameter vectors $\bm{\alpha}=(\alpha_{0},\ldots,\alpha_{K-1})\in\mathbb{C}^{K}$ for which $\mathcal{A}_{k}=\emptyset$ defines a proper affine variety in $\mathbb{C}^{K}$, and hence has Lebesgue measure zero. Therefore, for generic $\{\alpha_{k}:k\in[K]\}$, each $\mathcal{A}_{k}$ is a nonempty Zariski open subset of $\ker(\bm{U})$. Since $\ker(\bm{U})$ is a linear subspace and hence an irreducible affine variety, any finite intersection of nonempty Zariski open subsets of $\ker(\bm{U})$ is nonempty, i.e., $\mathcal{C}_{A}\neq\emptyset$. The details are provided in Appendix B. ∎
From the above discussion, for generic $\{\alpha_{k}:k\in[K]\}$, there exists a vector $\bm{c}\in\ker(\bm{U})$ such that the roots $\{\beta_{n}:n\in[N]\}$ of the polynomial $P(z)$ are disjoint from the given set $\{\alpha_{k}:k\in[K]\}$. This completes the proof of disjointness.
V-C3 Proof of Distinction
Recall that a polynomial $P(z)$ has $N$ distinct roots if and only if its discriminant $\mathrm{Disc}(P)\neq 0$. We first present the following lemma.
Lemma 4
Define a Zariski open subset of $\ker(\bm{U})$ as $\mathcal{C}_{D}\triangleq\{\bm{c}\in\ker(\bm{U}):\mathrm{Disc}(P)\neq 0\}$. For generic $\{\alpha_{k}:k\in[K]\}$, the set $\mathcal{C}_{D}$ is nonempty.
Proof:
The set of parameter vectors $\bm{\alpha}$ == $(\alpha_{0}$, $\ldots$, $\alpha_{K-1})$ $\in$ $\mathbb{C}^{K}$ for which $\mathcal{C}_{D}$ == $\emptyset$ defines a proper affine variety in $\mathbb{C}^{K}$, and therefore has Lebesgue measure zero. Hence, for generic $\{\alpha_{k}:k\in[K]\}$, we have $\mathcal{C}_{D}\neq\emptyset$. The details are provided in Appendix C. ∎
Combining Lemma 2, Lemma 3, and Lemma 4, we consider the intersection $\mathcal{C}_{N}\cap\mathcal{C}_{A}\cap\mathcal{C}_{D}$.
Lemma 5
For fixed $\{\alpha_{k}:k\in[K]\}$, if $\mathcal{C}_{N}$, $\mathcal{C}_{A}$, and $\mathcal{C}_{D}$ are nonempty, then $\mathcal{C}_{N}\cap\mathcal{C}_{A}\cap\mathcal{C}_{D}$ is a nonempty Zariski open and dense subset of $\ker(\bm{U})$.
Proof:
Since $\ker(\bm{U})$ is an irreducible affine variety, the intersection $\mathcal{C}_{N}\cap\mathcal{C}_{A}\cap\mathcal{C}_{D}$ of the nonempty Zariski open subsets $\mathcal{C}_{N}$, $\mathcal{C}_{A}$, and $\mathcal{C}_{D}$ is nonempty and Zariski open in $\ker(\bm{U})$. Moreover, any nonempty Zariski open subset of an irreducible affine variety is dense. Hence, $\mathcal{C}_{N}\cap\mathcal{C}_{A}\cap\mathcal{C}_{D}$ is dense in $\ker(\bm{U})$. This completes the proof of Lemma 5. ∎
From Lemma 5, for generic $\{\alpha_{k}:k\in[K]\}$, all but a Lebesgue measure-zero subset of vectors $\bm{c}\in\ker(\bm{U})$ satisfy the following properties: 1) $c_{N}\neq 0$, so that the normalized polynomial $P(z)=\frac{c_{0}}{c_{N}}+\frac{c_{1}}{c_{N}}z+\ldots+z^{N}$ is well defined; 2) $\bm{A}\bm{c}]_{k}\neq 0$ for all $k\in[K]$, ensuring that $\{\alpha_{k}:k\in[K]\}$ does not contain any root of $P(z)$; and 3) $\mathrm{Disc}(P)\neq 0$, so that $P(z)$ has $N$ distinct roots. This completes the proof of sufficiency of Theorem 2.
V-D Proof of Theorem 3
Theorem 3 follows directly from the constructive proof of sufficiency in Theorem 2.
VI Conclusion
This paper established the fundamental limits of CPA in distributed computing systems. Unlike existing polynomial coded computing schemes based on individual decoding, we showed that exact recovery of a weighted aggregation of polynomial evaluations can be achieved without reconstructing all individual polynomial evaluations. We derived a necessary and sufficient condition for the feasibility of CPA in the regime where individual decoding is infeasible. Based on this characterization, we determined the minimum number of responses required for feasible CPA for generic data points. Specifically, the minimum number of responses equals $\lfloor(K-1)/2\rfloor+1$ for linear computations and (d1)(K1)+1(d-1)(K-1)+1 for polynomial degree $d\geq 2$, both of which are strictly smaller than the d(K1)+1d(K-1)+1 responses required by individual decoding. In addition to the converse, we provided explicit constructions that achieve these limits.
Appendix A Proof of Lemma 2
We first show that $\mathcal{C}_{N}\neq\emptyset$ if and only if $\bm{u}_{N}\in\operatorname{span}\{\bm{u}_{0},\bm{u}_{1},\ldots,\bm{u}_{N-1}\}$, where $\bm{u}_{n}$ denotes the $n$-th column of $\bm{U}$. If $\mathcal{C}_{N}\neq\emptyset$, then there exists $\bm{c}=(c_{0},\dots,c_{N})^{\mathsf{T}}\in\ker(\bm{U})$ with $c_{N}\neq 0$. From $\bm{U}\bm{c}=\bm{0}$, we have $\sum_{n=0}^{N}c_{n}\bm{u}_{n}=\bm{0}$, which implies $c_{N}\bm{u}_{N}=-\sum_{n=0}^{N-1}c_{n}\bm{u}_{n}$. Since $c_{N}\neq 0$, dividing both sides by $c_{N}$ yields $\bm{u}_{N}=\sum_{n=0}^{N-1}\Big(-\frac{c_{n}}{c_{N}}\Big)\bm{u}_{n}$, and hence $\bm{u}_{N}\in\operatorname{span}\{\bm{u}_{0},\dots,\bm{u}_{N-1}\}$. Conversely, if $\bm{u}_{N}=\sum_{n=0}^{N-1}e_{n}\bm{u}_{n}$ for some $(e_{0},\dots,e_{N-1})\in\mathbb{C}^{N}$, we define $\bm{c}=(e_{0},\dots,e_{N-1},-1)^{\mathsf{T}}$. Then $\bm{U}\bm{c}=\sum_{n=0}^{N-1}e_{n}\bm{u}_{n}-\bm{u}_{N}=\bm{0}$, which implies $\bm{c}\in\ker(\bm{U})$ and $c_{N}=-1\neq 0$.
Hence, it suffices to show that $\bm{u}_{N}\in$ $\operatorname{span}\{\bm{u}_{0}$, $\bm{u}_{1}$, $\ldots$, $\bm{u}_{N-1}\}$ holds for generic $\bm{\alpha}\in\mathbb{C}^{K}$. From $N\geq C$, we consider the first $C$ columns of $\bm{U}$, denoted by $\bm{U}^{\prime}\triangleq$ $\bm{u}_{0}\ \bm{u}_{1}\ \ldots\ \bm{u}_{C-1}]\in\mathbb{C}^{C\times C}$. It suffices to show that $\operatorname{rank}(\bm{U}^{\prime})=C$, or equivalently $\det(\bm{U}^{\prime})\neq 0$, since this implies $\operatorname{span}\{\bm{u}_{0}$, $\dots$, $\bm{u}_{C-1}\}$ == $\mathbb{C}^{C}$, and hence $\bm{u}_{N}$ $\in$ $\operatorname{span}\{\bm{u}_{0}$, $\ldots$, $\bm{u}_{N-1}\}$. In the following, we will show that the collection of $\bm{\alpha}$ that leads to $\det(\bm{U}^{\prime})=0$ is a strict affine variety in $\mathbb{C}^{K}$, and hence has Lebesgue measure zero. It follows that for generic $\{\alpha_{k}:k\in[K]\}$, $\det(\bm{U}^{\prime})\neq 0$ and $\operatorname{rank}(\bm{U}^{\prime})=C$. Thus, $\bm{u}_{N}$ $\in$ $\operatorname{span}\{\bm{u}_{0}$, $\ldots$, $\bm{u}_{C-1}\}$ $\subseteq$ $\operatorname{span}\{\bm{u}_{0}$, $\ldots$, $\bm{u}_{N-1}\}$, and $\mathcal{C}_{N}\neq\emptyset$ for generic $\{\alpha_{k}:k\in[K]\}$.
In detail, we let $\bm{A}_{[:,\mathcal{S}]}$ denotes the sub-matrix of matrix $\bm{A}$ consisting of columns indexed by $\mathcal{S}$, and $\bm{A}_{[\mathcal{S},:]}$ denotes the sub-matrix consisting of rows indexed by $\mathcal{S}$. From $\bm{U}=\bm{V}\operatorname{diag}(\bm{w})\bm{A}$, we have $\bm{U}^{\prime}=\bm{V}\operatorname{diag}(\bm{w})\bm{A}_{[:,[C]]}$. From the definitions of $\bm{V}$ and $\bm{A}$, we have $\bm{V}=\bm{A}_{[:,[C]]}^{\mathsf{T}}$, and thus $\bm{U}^{\prime}=\bm{A}_{[:,[C]]}^{\mathsf{T}}\operatorname{diag}(\bm{w})\bm{A}_{[:,[C]]}.$ Applying the Cauchy–Binet formula to
$\bm{X}\triangleq\bm{A}_{[:,[C]]}^{\mathsf{T}}\operatorname{diag}(\bm{w})\in\mathbb{C}^{C\times K},\ \bm{Y}\triangleq\bm{A}_{[:,[C]]}\in\mathbb{C}^{K\times C},$
we obtain
$\det(\bm{U}^{\prime})=\sum_{\mathcal{S}\subseteq[K],\,|\mathcal{S}|=C}\det\!\big(\bm{X}_{[:,\mathcal{S}]}\big)\,\det\!\big(\bm{Y}_{[\mathcal{S},:]}\big).$ (16)
For each $\mathcal{S}\subseteq[K]$ with $|\mathcal{S}|=C$, we have $\bm{X}_{[:,\mathcal{S}]}$ == $\bm{A}_{[\mathcal{S},[C]]}^{\mathsf{T}}$ $\operatorname{diag}(\bm{w}_{\mathcal{S}})$ and $\bm{Y}_{[\mathcal{S},:]}$ == $\bm{A}_{[\mathcal{S},[C]]}$, where $\bm{w}_{\mathcal{S}}\triangleq(w_{k})_{k\in\mathcal{S}}$. Therefore, $\det\!\big(\bm{X}_{[:,\mathcal{S}]}\big)=\det(\bm{A}_{[\mathcal{S},[C]]}^{\mathsf{T}})\det(\operatorname{diag}(\bm{w}_{\mathcal{S}}))=\Big(\prod_{k\in\mathcal{S}}w_{k}\Big)\det(\bm{A}_{[\mathcal{S},[C]]})$, and (16) becomes
$\det(\bm{U}^{\prime})=\sum_{\mathcal{S}\subseteq[K],\,|\mathcal{S}|=C}\Big(\prod_{k\in\mathcal{S}}w_{k}\Big)\det\!\big(\bm{A}_{[\mathcal{S},[C]]}\big)^{2}.$
Each $\bm{A}_{[\mathcal{S},[C]]}$ is a $C\times C$ Vandermonde matrix built from $\{\alpha_{k}:k\in\mathcal{S}\}$, and hence its determinant is a nonzero polynomial whenever the $\alpha_{k}$’s are distinct. Consequently, $\det(\bm{U}^{\prime})$ is a nonzero polynomial in $\{\alpha_{k}:k\in[K]\}$. Therefore, the set $\{\bm{\alpha}\in\mathbb{C}^{K}:$ $\det(\bm{U}^{\prime})=0\}$ is a strict affine variety in $\mathbb{C}^{K}$. From the discussion above, we conclude that $\mathcal{C}_{N}\neq\emptyset$ for generic $\{\alpha_{k}:k\in[K]\}$.
Appendix B Proof of Lemma 3
B-A $\mathcal{A}_{k}\neq\emptyset$ for Generic $\{\alpha_{k}:k\in[K]\}$
Assume that, for contradiction,
$\bm{A}\bm{c}]_{k}=0\quad\text{for all }\bm{c}\in\ker(\bm{U}).$ (17)
Then the $k$-th row of $\bm{A}$, denoted by $\bm{a}_{k}$, lies in the row space of $\bm{U}$, i.e., $\bm{a}_{k}\in\operatorname{span}\{\bm{u}^{0},\bm{u}^{1},\ldots,\bm{u}^{C-1}\}$, where $\bm{u}^{i}$ denotes the $i$-th row of $\bm{U}$. We consider two cases depending on $d$.
B-A1 When $d\geq 2$
From assumption, there exist coefficients $\lambda_{0},\ldots,\lambda_{C-1}$ such that $\bm{a}_{k}$ $\sum_{i=0}^{C-1}\lambda_{i}\bm{u}^{i}$. Using the factorization $\bm{U}$ == $\bm{V}$ $\operatorname{diag}(\bm{w})$ $\bm{A}$, let $\bm{v}^{i}$ denote the $i$-th row of $\bm{V}$, whose $l$-th entry is $v^{i}[l]=\alpha_{l}^{\,i}$. Then $\bm{u}^{i}$ == $\bm{v}^{i}$ $\operatorname{diag}(\bm{w})$ $\bm{A}$ == $\sum_{l=0}^{K-1}\alpha_{l}^{\,i}w_{l}\bm{a}_{l}$. Substituting this into the expression for $\bm{a}_{k}$ and exchanging the order of summation yields $\bm{a}_{k}=\sum_{l=0}^{K-1}w_{l}\Big(\sum_{i=0}^{C-1}\lambda_{i}\alpha_{l}^{\,i}\Big)\bm{a}_{l}$. Define the degree-(C1)(C-1) polynomial $Q(z)\triangleq\sum_{i=0}^{C-1}\lambda_{i}z^{i}$. Then $\bm{a}_{k}=\sum_{l=0}^{K-1}w_{l}Q(\alpha_{l})\bm{a}_{l}$. Let $x_{l}\triangleq w_{l}Q(\alpha_{l})$ for $l\in[K]$. Comparing the $n$-th coordinate of $\bm{a}_{k}$ gives
$\sum_{l=0}^{K-1}x_{l}\alpha_{l}^{n}=\alpha_{k}^{n},\qquad n\in[N+1].$ (18)
Under the sufficiency condition of Theorem 2 for $d\geq 2$, i.e., $N\geq K$, we restricting (18) to $n=0,1,\ldots,K-1$ yields the Vandermonde system $\bm{V}_{K}\bm{x}=\bm{b}_{k}$, where $(\bm{V}_{K})_{n,l}=\alpha_{l}^{n}$ and $\bm{b}_{k}=(\alpha_{k}^{0},\alpha_{k}^{1},\ldots,\alpha_{k}^{K-1})^{\mathsf{T}}$ is the $k$-th column of $\bm{V}_{K}$. Since the $\{\alpha_{l}\}$ are distinct, the Vandermonde matrix $\bm{V}_{K}$ is invertible. Moreover, $\bm{b}_{k}$ is exactly the $k$-th column of $\bm{V}_{K}$. Hence the unique solution to $\bm{V}_{K}\bm{x}=\bm{b}_{k}$ is $\bm{x}=\bm{e}_{k}$. Equivalently, $w_{l}Q(\alpha_{l})=0$ for all $l\neq k$, and $w_{k}Q(\alpha_{k})=1$. Thus $Q(\alpha_{l})=0$ for all $l\neq k$. However, $\deg(Q)\leq C-1$ and $C<K$, so $Q$ cannot have K1K-1 distinct zeros for generic distinct $\{\alpha_{l}\}$. This contradiction shows that $\bm{a}_{k}\notin\operatorname{span}\{\bm{u}^{0},\bm{u}^{1},\ldots,\bm{u}^{C-1}\}$ for generic $\{\alpha_{k}:k\in[K]\}$.
B-A2 When $d=1$
Condition (17) is equivalent to $\bm{a}_{k}\bm{c}=0$ for all $\bm{c}\in\ker(\bm{U})$. This implies
$\bm{a}_{k}\in(\ker(\bm{U}))^{\perp}=\operatorname{col}(\bm{U}^{\mathsf{T}}).$ (19)
Next, we show that (19) cannot hold for generic $\{\alpha_{k}:k\in[K]\}$. Since $\bm{U}\in\mathbb{C}^{C\times(N+1)}$, we have $\dim\operatorname{col}(\bm{U}^{\mathsf{T}})$ == $\operatorname{rank}(\bm{U})$ $\leq$ $C<N+1$. On the other hand, $\bm{a}_{k}=$ $(1$, $\alpha_{k}$, $\alpha_{k}^{2}$, $\ldots$, $\alpha_{k}^{N})$ depends on $\alpha_{k}$. The condition $\bm{a}_{k}\in\operatorname{col}(\bm{U}^{\mathsf{T}})$ therefore imposes a nontrivial algebraic constraint on $\{\alpha_{k}:k\in[K]\}$, and hence holds only on a strict affine variety of $\mathbb{C}^{K}$. Consequently, for generic $\{\alpha_{k}:k\in[K]\}$, (19) does not hold, which contradicts (17).
Combining the two cases, we conclude that for generic $\{\alpha_{k}:k\in[K]\}$, there exists $\bm{c}\in\ker(\bm{U})$ such that $\bm{A}\bm{c}]_{k}\neq 0$. Hence $\mathcal{A}_{k}\neq\emptyset$.
B-B $C_{A}\neq\emptyset$ for Generic $\{\alpha_{k}:k\in[K]\}$
Since $\ker(\bm{U})$ is a linear subspace and hence an irreducible affine variety, any finite intersection of nonempty Zariski open subsets of $\ker(\bm{U})$ is nonempty. Hence, $\mathcal{C}_{A}\neq\emptyset$.
Appendix C Proof of Lemma 4
Since $P(z)=P_{\mathrm{form}}(z)/c_{N}$ with $c_{N}\neq 0$, the discriminants $\mathrm{Disc}(P)$ and $\mathrm{Disc}(P_{\mathrm{form}})$ differ only by a nonzero multiplicative constant. Hence, $\mathrm{Disc}(P)\neq 0$ if and only if $\mathrm{Disc}(P_{\mathrm{form}})\neq 0$. Recall that $P_{\mathrm{form}}(z)=\sum_{n=0}^{N}c_{n}z^{n}$ and that $\mathrm{Disc}(P_{\mathrm{form}})$ is a polynomial in the coefficients $\bm{c}$. Moreover, $\mathrm{Disc}(P_{\mathrm{form}})$ is not identically zero, since there exist polynomials with distinct roots. Hence, $\mathcal{O}\triangleq\{\bm{c}\in\mathbb{C}^{N+1}:\mathrm{Disc}(P_{\mathrm{form}})\neq 0\}$ is a nonempty Zariski open subset of $\mathbb{C}^{N+1}$.
Fix $\{\alpha_{k}:k\in[K]\}$ and the corresponding matrix $\bm{U}$. The set $\mathcal{C}_{D}=\{\bm{c}\in\ker(\bm{U}):\mathrm{Disc}(P_{\mathrm{form}})\neq 0\}$ is the intersection of $\ker(\bm{U})$ with the Zariski open set $\mathcal{O}$, and is therefore Zariski open in $\ker(\bm{U})$. Moreover, $\mathcal{C}_{D}$ is empty if and only if $\mathrm{Disc}(P_{\mathrm{form}})$ vanishes identically on $\ker(\bm{U})$.
Since the entries of $\bm{U}$ depend polynomially on $\bm{\alpha}$, the set
$\mathcal{B}$ $\triangleq\{\bm{\alpha}\in\mathbb{C}^{K}:\mathcal{C}_{D}=\emptyset\}$
$=\{\bm{\alpha}:\mathrm{Disc}(P_{\mathrm{form}})\text{ vanishes identically on }\ker(\bm{U}(\bm{\alpha}))\}$
is an affine variety of $\mathbb{C}^{K}$. Moreover, $\mathcal{B}$ is proper because there exists at least one choice of $\bm{\alpha}$ for which $\mathcal{C}_{D}\neq\emptyset$. Consequently, for generic $\{\alpha_{k}:k\in[K]\}$, the set $\mathcal{C}_{D}$ is nonempty.